OpenStudy (anonymous):

you have a bag with 20 quarters, 10 dimes, 30 nickels and 15 pennies. If you reach into the bag and randomly draw out a coin what is the probability that its not a dime?

OpenStudy (anonymous):

Please help me!!!

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Two questions : how many coins are there in total ? how many coins are NOT dimes ?

OpenStudy (anonymous):

there are 75 coins total and 65 coins are not dimes

ganeshie8 (ganeshie8):

yes ! probability for "\(\large \text{not dime}\)" = \(\large \dfrac{ \text{ # of not dimes} }{ \text{total #}}\)

ganeshie8 (ganeshie8):

just take the ratio

OpenStudy (anonymous):

But what would the exact percent be?

OpenStudy (paki):

65/75 = 13/15

ganeshie8 (ganeshie8):

probability is always between 0 and 1 to get the percentage multiply the ratio by 100